OpenStudy (anonymous):

a bullet is moving with a speed of 367m/s when it enters a lump of moist clay. the bullet penetrates a distance of 0.0621m. determine the acceleration of the bullet while moving into the clay. assume uniform acceleration.

OpenStudy (anonymous):

\[v ^{2}=u ^{2} +2ad\]Let, v = final velocity (0) u = intial velocity (367m/s) d = 0.0621 Solve for a, \[0^{2}=367^{2}+2\times a \times 0.0621\]